OpenStudy (anonymous):

solve for M: 2/3 = m/24 i got m=28

OpenStudy (anonymous):

2/3=m/24 2/3*24=m m=48/3=16

OpenStudy (zzr0ck3r):

\[\frac{2}{3}=\frac{m}{24}\\m=24*\frac{2}{3}\]